OpenStudy (anonymous):

HELP AWARDING METAL!! Verify each trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation. (sin x)(tan x cos x - cot x cos x) = 1 - 2 cos2x

OpenStudy (jdoe0001):

soooo \(\bf sin(x)\left[ tan(x)cos(x)-cot(x)cos(x) \right]=1-2cos^2(x) \\ \quad \\ \quad \\ sin(x)\left[ {\color{brown}{ \cfrac{sin(x)}{\cancel{cos(x)}} }} \cancel{cos(x)}-{\color{brown}{ \cfrac{cos(x)}{sin(x)}}} cos(x) \right] \\ \quad \\ sin(x)\left[ sin(x)-\cfrac{cos^2(x)}{sin(x)} \right]\implies sin(x)sin(x)-\cancel{sin(x)}\cfrac{cos^2(x)}{\cancel{sin(x)}}\) what would you end up with there?